OpenStudy (anonymous):

Determine whether the given linear equations are parallel, perpendicular, or neither. y = -9x +11

OpenStudy (anonymous):

your question requires two equations or more, you have only supplied one!

OpenStudy (anonymous):

I think its y=1/9x + 8

OpenStudy (anonymous):

ah okay so linear means line. If you're talking parallel the lines have a slope (m) that is equal. If the lines are perpendicular you have a slope that is the negative reciprocal. and both of your equations are in the form y = mx+b

OpenStudy (anonymous):

your slope for equation one is -9, your slope for equation two is +1/9 making it perpendicular.

OpenStudy (anonymous):

okay. so would this one 8x - 4y = - 12 8x - 4y = 8 be parallel?

OpenStudy (anonymous):

for those you would have to put them in y = mx+b form first, so you would need to solve for Y and then evaluate if the "m" or the "slope" is the same.

OpenStudy (anonymous):

i just solved both for Y on paper and according to my arithmetic they are not parallel, they have different slopes. You should check it on your own though.

OpenStudy (anonymous):

They are parallel. The only difference is the constant. Putting both of the equations in slope intercept form results in a slope of 2 for each equation, making them parallel.

OpenStudy (anonymous):

calmat01 is correct, i redid the problems they are parallel, sorry!

OpenStudy (anonymous):

alright. ty